# CV-VAE: A Compatible Video VAE for Latent Generative Video Models

### v1

@@ 4.2 Image and Video Reconstruction
As illustrated in Tab. 1, we present the parameter count (Params), Frame Compression Ratio (FCR), and compatibility with existing diffusion models (Comp.) for various VAE models. Thanks to the latent constraint, our model is compatible with current diffusion models, compresses videos by 4$\times$ in the temporal dimension, and achieves top-tier image and video reconstruction quality. This enables the generation of longer videos under roughly the same computational resources.
We also conducted a qualitative comparison of the reconstruction results for different VAE models, as shown in Fig. 4. In the top row, we reconstructed images with a resolution of $512\times 512$ and compared them with Image VAE models. All three models compressed the images to a latent size of $64\times 64$. Our results were close to those of VAE-SD2.1, while VQGAN had the worst performance. In the bottom row, we reconstructed videos with a resolution of $33\times 512\times 512$ and compared them with Video VAE models. All three models compressed the videos to a latent size of $9\times 64\times 64$. Comparing the decoded videos at the same frames, our model achieved the best results. Check Appendx A.2 and A.3 for more reconstruction results.
@@ 4.3 Compatibility with Existing Models
Image Models
@@
Trainable COCO2017-Val
FID($\downarrow$) CLIP($\downarrow$) PIC($\uparrow$)
SD2.1 [26] $\times$ 57.3 0.312 0.354
SD2.1+CV-VAE $\times$ 57.6 0.311 0.360
Table 2: Quantitative results of text-to-image generation.
Video Models
Method Trainable FCR Frames UCF-101 MSR-VTT
SVD [2] $\times$ 1x 25 402 8.20 0.791 310 1.30 0.588
SVD+CV-VAE-I $\times$ 1x 25 419 6.73 0.763 262 1.67 0.609
@@
SVD+CV-VAE-V Output layer 4x 97 681 13.1 0.858 295 2.26 0.734
@@
As shown in Tab. 3, incorporating CV-VAE-I into a frozen SVD immediately yields video generation quality comparable to the original VAE. Using CV-VAE in video mode can also decode videos generated by SVD, and further improvements in video decoding quality can be achieved by fine-tuning only the output layer (approximately 12k parameters). One of the reasons for the noticeable gap in test metrics between SVD+CV-VAE-I and SVD+CV-VAE-V is that they use different numbers of frames, making a direct comparison challenging.
4.4 Ablation Study
Influence of Regularization Type
@@ A.1 CV-VAE Model Architecture
A.2 Qualitative Examples of Image Reconstruction
@@
A.3 Qualitative Examples of Video Reconstruction
@@
A.4 Compatibility with Existing Text-to-video Model

### v2

@@ 4.2 Image and Video Reconstruction
As illustrated in Tab. 1, we present the parameter count (Params), Frame Compression Ratio (FCR), and compatibility with existing diffusion models (Comp.) for various VAE models. Thanks to the latent constraint, our model is compatible with current diffusion models, compresses videos by 4$\times$ in the temporal dimension, and achieves top-tier image and video reconstruction quality. This enables the generation of longer videos under roughly the same computational resources. Reconstruction quality improves as the number of latent channels increases. For comparison results with 16 latent channels, please refer to Appendx A.2.
We also conducted a qualitative comparison of the reconstruction results for different VAE models, as shown in Fig. 4. In the top row, we reconstructed images with a resolution of $512\times 512$ and compared them with Image VAE models. All three models compressed the images to a latent size of $64\times 64$. Our results were close to those of VAE-SD2.1, while VQGAN had the worst performance. In the bottom row, we reconstructed videos with a resolution of $33\times 512\times 512$ and compared them with Video VAE models. All three models compressed the videos to a latent size of $9\times 64\times 64$. Comparing the decoded videos at the same frames, our model achieved the best results. Check Appendx A.3 and A.4 for more reconstruction results.
@@ 4.3 Compatibility with Existing Models
Trainable COCO2017-Val
FID($\downarrow$) CLIP($\downarrow$) PIC($\uparrow$)
SD2.1 [28] $\times$ 57.3 0.312 0.354
SD2.1+CV-VAE $\times$ 57.6 0.311 0.360
Table 2: Quantitative results of text-to-image generation.
Text-to-Image Models
@@
Image-to-Video Models
Method Trainable FCR Frames UCF-101 MSR-VTT
SVD [4] $\times$ 1$\times$ 25 402 8.20 0.791 310 1.30 0.588
SVD+CV-VAE-I $\times$ 1$\times$ 25 419 6.73 0.763 262 1.67 0.609
[…]
SVD+CV-VAE-V Output layer 4$\times$ 97 681 13.1 0.858 295 2.26 0.734
@@
As shown in Tab. 3, incorporating ‘CV-VAE-I’ into a frozen SVD immediately yields video generation quality comparable to the original VAE. Using CV-VAE in video mode can also decode videos generated by SVD, and further improvements in video decoding quality can be achieved by fine-tuning only the output layer (approximately 12k parameters). One of the reasons for the noticeable gap in test metrics between ‘SVD+CV-VAE-I’ and ‘SVD+CV-VAE-V’ is that they use different numbers of frames, making a direct comparison challenging.
      Method       FVD($\downarrow$)       KVD($\downarrow$)       PIC($\uparrow$)
      SVD+RIFE [18]       253       3.12       0.721
      SVD+CV-VAE       295       2.26       0.734
Table 4: Comparison between CV-VAE and frame interpolation model.
By fine-tuning a small number of parameters, the image-to-video model can generate smoother and longer videos through CV-VAE, effectively serving as a frame interpolation method. Therefore, we compared CV-VAE with existing video interpolation models [18], conducting experiments on MSR-VTT [39]. We first used SVD to generate a video of size $25\times 576\times 1024$, and then applied the interpolation model to expand the video from 25 frames to 97 frames. For our results, we directly generated a video of size $97\times 576\times 1024$ using CV-VAE and the fine-tuned SVD. The comparison results are shown in Tab. 4, where CV-VAE outperformed the FIRE [18] in two out of three metric, validating the potential of CV-VAE as an interpolation model.
Text-to-Video Models
In this section, we integrate CV-VAE into the existing text-to-video model to futher validate the effectiveness. ‘VC2’ refers to decoding the results generated by VideoCrafter2 [7] using the original 2D VAE, ‘VC2+CV-VAE-I’ indicates decoding the results using CV-VAE in image mode, and ‘VC2+CV-VAE-V’ denotes decoding the results using CV-VAE in video mode, which generates videos of 4×4$\times$4 × frames. We only fine-tuned a small number of parameters in U-Net, including the first and last layers. We use captions from the validation set of MSR-VTT [39] for evaluation, with the resolution of 320×512. Following the approach taken by previous studies [37], we used the CLIP [26] metric to evaluate the generation quality of text-to-video models, including Frame Consistency (F.C.) and Textual Alignment (T.A.). The experimental results are shown in Tab. 5, where the ‘VC2+CV-VAE-V’ setting achieved the best generation performance through fine-tuning VideoCrafter2. Check Appendx A.5 for quantitative comparison.
Methods FCR Trainable CLIP Metric
F.C. T.A.
VC2 [7] 1$\times$ $\times$ 0.292 0.960
VC2+CV-VAE-I 1$\times$ $\times$ 0.290 0.964
@@
VC2+CV-VAE-V 4$\times$ $\checkmark$ 0.301 0.987
Table 5: Evaluation results of text-to-video generation.
4.4 Ablation Study
@@ A.1 CV-VAE Model Architecture
A.2 CV-VAE with more latent channels
Method Params FCR Comp. COCO-Val Webvid-Val
VAE-SD3 [13] 34M + 49M 1x - 32.6 0.899 0.064 33.4 0.910 0.064
CVVAE-SD3 68M + 114M 4x $\checkmark$ 33.6 0.916 0.072 32.0 0.899 0.087
Table 8: Quantitative evaluation on image and video reconstruction between. FCR represents the frame compression rate, and Comp. indicates compatibility with existing generative models.
More latent channels generally lead to better reconstruction performance in VAEs [13, 10], which is crucial for image and video editing tasks. In this subsection, I conducted experiments using a VAE with more latent channels (z=16). We used the 2D VAE from SD3 [13] as the baseline and trained the CVVAE-SD3 based on latent regularization. Testing was conducted under the same settings as in Tab. 1. The comparison results are shown in Tab. 8. CVVAE-SD3 outperforms VAE-SD3 in image reconstruction but is inferior to VAE-SD3 in video reconstruction. The main reason is that CVVAE-SD3 has a higher compression ratio in video compression, resulting in the loss of more information, though its reconstruction quality is significantly higher than that of VAE with z=4.
A.3 Qualitative Examples of Image Reconstruction
@@
A.4 Qualitative Examples of Video Reconstruction
@@
A.5 Compatibility with Existing Text-to-video Model